BADKARMA88:

Can someone help me with my Math. It is -6y+4=-8

martai:

MK I got u gimme like 2 min

martai:

dont give out the answer

anthony5012005:

so u subtact 4 from both sides then divide -6

BADKARMA88:

Mk

martai:

Subtract 4 from both sides -6y+4=-8 6y+4-4=8-4 simplify and subtract -6y=12

martai:

then u wanna divide both sides by the same factor

BADKARMA88:

I got it ty

martai:

yw(:

martai:

then u get ur answer

BADKARMA88:

yea

3lout:

ok thx

martai:

which is ?

BADKARMA88:

y=2

3lout:

y=2

martai:

yup good job

JordynTheAlien:

You need to simplify. Then divide by both sides, then simplify again.

3lout:

@jordynthealien wrote:
You need to simplify. Then divide by both sides, then simplify again.
ight